OpenStudy (anonymous):

Show that the equation represents a circle by rewriting it in standard form, and find the center and radius of the circle. x^2+y^2-1/4x+1/4y=7/32

OpenStudy (tkhunny):

Well, do it. You will need to "Complete the Square", twice. x^2 - (1/4)x + _____ + y^2 + (1/4)y + ____ = (7/32) + _____ + ____ Go!

OpenStudy (anonymous):

I need help

OpenStudy (tkhunny):

You're saying you do not know how to "Complete the Square"?

OpenStudy (anonymous):

no i do

OpenStudy (anonymous):

let me do it

OpenStudy (tkhunny):

Excellent. Let's see what you get. I rearranged it a little more conveniently for you up above.

OpenStudy (anonymous):

How do I not put it in decimal form?

OpenStudy (tkhunny):

Simply don't. x^2 - (1/4)x + _____ + y^2 + (1/4)y + ____ = (7/32) + _____ + ____ (1/2)*(1/4) = 1/8 (1/8)^2 = 1/64 This gives: x^2 - (1/4)x + (1/64) + y^2 + (1/4)y + ____ = (7/32) + (1/64) + ____ How about the other one?

OpenStudy (anonymous):

y^2+1/4y+1/64

OpenStudy (tkhunny):

Okay, now we have x^2 - (1/4)x + (1/64) + y^2 + (1/4)y + (1/64) = (7/32) + (1/64) + (1/64) What's next? There should be some significant simplification or we just failed to make any sense at all.

OpenStudy (anonymous):

for the right side I go =1/4

OpenStudy (tkhunny):

7/32 + 1/64 + 1/64 = 7/32 + 2/64 = 7/32 + 1/32 = 8/32 = 1/4 Okay, I'll bite. How about the left-hand side?

OpenStudy (anonymous):

I am not quite sure

OpenStudy (tkhunny):

What was the point of "Completing the Square"?

OpenStudy (anonymous):

To get the factors

OpenStudy (tkhunny):

No, it was to make perfect squares! y^2+1/4y+1/64 = (y + 1/8)^2 Do that with the x and write the whole thing. You'll be almost done.

OpenStudy (anonymous):

so it would be (x-1/8)^2

OpenStudy (tkhunny):

Yes. Now, we are ready to answer all the Problem Statement asks. (x - 1/8)^2 + (y + 1/8)^2 = 1/4 This certainly looks like a circle to me. What is the Center? What is the Radius?

OpenStudy (anonymous):

center= 1/8, -1/8

OpenStudy (anonymous):

radius= (1/4)

OpenStudy (anonymous):

wait or is the radius the square root of 1/4?

OpenStudy (tkhunny):

You got it!!! I deliberately failed to write it in the proper form, just to see if you were paying attention. (x - 1/8)^2 + (y + 1/8)^2 = (1/2)^2 Now, it's a bit more obvious.

OpenStudy (anonymous):

YAY! Thank you so much!!

OpenStudy (tkhunny):

Good work. Recognizing the standard form was the important part of the problem. Completing the square was the important process to get us there.